OpenStudy (anonymous):

if the ratio of 4 to b is 5 to 7 then b=?

jimthompson5910 (jim_thompson5910):

"if the ratio of 4 to b is 5 to 7" translates to 4/b = 5/7 Solve this for b 4/b = 5/7 4*7 = 5b 28 = 5b 28/5 = b b = 28/5